Problem 5-7 (LG 5-2) You can purchase a T-bill that is 90 days from maturity for $10,465. The T-bill has a face value of $10,500. a. Calculate the T-bill’s quoted yield. (Use 360 days in a year. Do not round intermediate calculations. Round your answer to 3 decimal places. (e.g., 32.161)) b. Calculate the T-bill’s bond equivalent yield. (Use 365 days in a year. Do not round intermediate calculations. Round your answer to 3 decimal places. (e.g., 32.161)) c. Calculate the T-bill’s EAR. (Use 365 days in a year. Do not round intermediate calculations. Round your answer to 3 decimal places. (e.g., 32.161))
Answer #1

a. T-bill’s quoted yield = [($10,500 - $10,465) / $10,500] × (360/90) = 1.333%

b. T-bill’s bond equivalent yield = [($10,500 - $10,465) / $10,465] × (365/90) = 1.356%

c. T-bill’s EAR = [1 + 0.01356/(365/90)](365/90) − 1 = 1.363%